# Partial identification of principal causal effects under violations of principal ignorability

### v1

@@ Abstract
Principal stratification is a general framework for studying causal mechanisms involving post-treatment variables. When estimating principal causal effects, the principal ignorability assumption is commonly invoked, which we study in detail in this manuscript. Our first key contribution is studying a commonly used strategy of using parametric models to jointly model the outcome and principal strata without requiring the principal ignorability assumption. We show that even if the joint distribution of principal strata is known, this strategy necessarily leads to only partial identification of causal effects, even under very simple and correctly specified outcome models. While principal ignorability can lead to point identification in this setting, we discuss alternative, weaker assumptions and show how they lead to more informative partial identification regions. An additional contribution is that we provide theoretical support to strategies used in the literature for identifying association parameters that govern the joint distribution of principal strata. We prove that this is possible, but only if the principal ignorability assumption is violated. Additionally, due to partial identifiability of causal effects even when these association parameters are known, we show that these association parameters are only identifiable under strong parametric constraints. Lastly, we extend these results to more flexible semiparametric and nonparametric Bayesian models.
@@ 2.1 Notation and estimands
Throughout let $\boldsymbol{X}$ denote a $p$-dimensional vector of pre-treatment covariates, $T$ a binary treatment, $S$ a binary or continuous intermediate, and $Y$ an outcome of interest. Our observed data consists of $n$ independent copies of these random variables, and we refer to the observed data by $\mathcal{D}$. We adopt the potential outcome framework [34] and define $S(t)$ to be the potential intermediate that would be observed had treatment been set to $t$, and similarly let $Y(t)$ be the potential outcome. Throughout, we let $\boldsymbol{U}=(S(0),S(1))$ denote the joint values of the two potential intermediates, and let $S$ be the observed value of the intermediate. We assume the Stable Unit Treatment Value Assumption (SUTVA) [35], which implies that 1) there is no ’interference,’ meaning that the potential outcomes for unit $i$ do not change with the treatments assigned to other units, and 2) there are not different versions of each treatment level. This assumption ensures that $S=S(T)$ and $Y=Y(T)$. Principal strata are defined by the joint potential values of the intermediate variable $\boldsymbol{U}=(S(0),S(1))$. The causal estimands of interest are then defined in terms of average differences in potential outcomes within principal strata, which are given by
$PCE(\boldsymbol{u})=E[Y(1)-Y(0)\mid\boldsymbol{U}=\boldsymbol{u}].$
@@
$PCE(\boldsymbol{u},\boldsymbol{x})=E[Y(1)-Y(0)\mid\boldsymbol{U}=\boldsymbol{u},\boldsymbol{X}=\boldsymbol{x}].$
@@ 2.2 Review of identification assumptions in PS analysis
0 ¡ $P(T=1\mid\boldsymbol{X}=\boldsymbol{x})<1$ for all $\boldsymbol{x}$.
@@ 2.2.1 Binary intermediates
This assumption rules out a direct effect of the treatment on the outcome that does not go through a change in the intermediate. Again, this is reasonable in randomized studies of noncompliance where randomization to a treatment should not affect the outcome, except through its impact on what treatment is actually taken. Under the exclusion restriction and monotonicity assumptions, the average causal effect for compilers is identified as the ratio of the average causal effects on $Y$ and $S$:
$E[Y(1)-Y(0)\mid S(0)=0,S(1)=1]=\frac{E[Y(1)-Y(0)]}{E[S(1)-S(0)]}.$
@@
$E[Y(t)\mid S(t),S(1-t),\boldsymbol{X}]=E[Y(t)\mid S(t),\boldsymbol{X}],\qquad t=0,1.$
@@
$PCE(\boldsymbol{u},\boldsymbol{x})=E[Y\mid S=s_{1},\boldsymbol{X}=\boldsymbol{x},T=1]-E[Y\mid S=s_{0},\boldsymbol{X}=\boldsymbol{x},T=0],$
@@ 2.2.2 Continuous intermediates
For a binary $S$, there are only four principal strata, and the monotonicity assumption eliminates one of these groups and enables the identification of the probability mass of $\boldsymbol{U}\mid\boldsymbol{X}$ ([1]). For a continuous $S$, there are infinitely many principal strata, and while the monotonicity assumption can rule out some of them, there still remain infinitely many principal strata. Thus, monotonicity alone cannot produce the identification of the distribution of $\boldsymbol{U}\mid\boldsymbol{X}$, sometimes referred to as the principal density for continuous intermediates. Typically, researchers make parametric assumptions in order to identify this distribution. One approach is to assume the joint distribution of principal strata follows a known copula function $\mathbb{C}_{\rho}(\cdot,\cdot)$ (see [28]; [20]; [4])
$P(S(0)\leq s_{0},S(1)\leq s_{1}\mid\boldsymbol{X})=\mathbb{C}_{\rho}(P(S(0)\leq s_{0}\mid\boldsymbol{X}),P(S(1)\leq s_{1}\mid\boldsymbol{X})),$
@@
In general, the explicit identification of the correlation between potential intermediates is infeasible because only one of them is ever observed per observation. This has led to a common practice of $\rho$ being selected based on domain knowledge, and a sensitivity analysis is performed to assess the impact of this parameter. Throughout this article, we refer to parameters like $\rho$ that cannot be identified marginally as the association parameters of the principal strata. Under this distributional assumption, combined with Assumptions (1) and (5), [28] demonstrated nonparametric identification of PCEs. Given that knowledge of this correlation parameter is a strong assumption, and that results can be sensitive to its choice, many authors have taken a parametric modeling approach to estimate this correlation. Additionally, in most of these works, principal ignorability is not assumed, showing that the parametric modeling assumptions reduce the number of structural assumptions that need to be made. This is the focus of our work, as we describe the extent to which inference in these settings is possible in the following sections.
@@ 3 On the possibility of identifying $\rho$
3.1 A simple Bayesian parametric model
@@
For simplicity of discussion and derivations, we first focus on a simple, linear parametric model for the outcome as well as the principal strata. We extend these ideas to semiparametric and nonparametric models in Section 7. We focus on a model given by
$\begin{split}E[Y_{i}\mid\boldsymbol{U}_{i},\boldsymbol{X}_{i},T_{i}=t]&=\boldsymbol{\beta}_{t}^{T}\boldsymbol{U}_{i}+\lambda_{t}+\boldsymbol{\gamma}^{T}\boldsymbol{X}_{i},\\ E[\boldsymbol{U}_{i}\mid\boldsymbol{X}_{i}]&=(\phi_{0}+\boldsymbol{\alpha}^{T}\boldsymbol{X}_{i},\phi_{1}+\boldsymbol{\alpha}^{T}\boldsymbol{X}_{i})^{T},\end{split}$ (1)
where $\boldsymbol{\beta}_{t}=(\beta_{t0},\beta_{t1})$ for $t=0$, 1. It may appear that we are focusing on overly simple models, but in what follows, we will see that even in such simple, parametric situations, issues of model identifiability arise, which will have implications for more complex model structures.
As mentioned earlier, we primarily focus on settings with a continuous intermediate variable, though analogous results hold for the binary intermediate variable setting. We assume Gaussian errors in the outcome model and a joint Gaussian distribution for the principal strata. We define $\mu_{yit}$ to be the conditional mean of the potential outcomes, $E[Y_{it}\mid\boldsymbol{U}_{i},\boldsymbol{X}_{i},T_{i}=t]$, and $\boldsymbol{\mu}_{si}=(\mu_{si0},\mu_{si1})$ to be the conditional mean of principal strata. Thus, a full data generating model can be written as:
@@
In model (2), $\rho$ is the association parameter of the principal strata and plays an important role in estimating PCEs. To the best of our knowledge, there is no literature formally studying whether it is possible to identify the association parameters, which we do in the following section.
@@ 3.2 Identification of the association parameter of the principal strata
For a continuous intermediate, estimating average PCEs always requires estimating the principal density, which involves the association parameters. Identifying the association parameters is generally infeasible due to the lack of joint observations of principal strata. Some researchers assume these are known ([21], while others have tried to estimate them using similar models to those in 2 ([37], [4]). The idea is that although $S(0)$ and $S(1)$ are never jointly observed, their association is implicitly included in the outcome model. The following result provides theoretical justification to this strategy:
@@
Suppose that $\boldsymbol{\theta}=(\boldsymbol{\beta}_{0},\boldsymbol{\beta}_{1},\lambda_{0},\lambda_{1},\sigma_{y}^{2},\boldsymbol{\gamma},\boldsymbol{\alpha},\sigma_{s0},\sigma_{s1})$ are known and principal ignorability fails: that is, at least one of $\beta_{01}$ or $\beta_{10}$ is nonzero. Under model 2, the posterior mode of $\rho\mid\mathcal{D}$ is consistent.
Theorem 1 shows when principal ignorability is violated, then identification of $\rho$ is indeed possible, though this result should be interpreted with caution. For one, it required the outcome model to be correctly specified with known parameters. We will see in the subsequent sections that the model parameters are only partially identified in many settings, and the implications of Theorem 1 would only be useful for estimating $\rho$ if we can consistently estimate these regression parameters. Another issue is that this result relies on principal ignorability being violated. One can show that the posterior of $\rho$ reduces to the prior distribution for $\rho$ when principal ignorability holds. This can also be seen in Theorem 2, where we show the asymptotic variance of the posterior distribution of $\rho$.
@@
$\begin{split}Var(\rho\mid\mathcal{D})&\approx n^{-1}\bigg{(}\bar{T}{\beta_{10}^{\ast}}^{2}{\sigma_{s0}^{\ast}}^{2}\bigg{(}\frac{2{{\rho}^{\ast}}^{2}{\beta_{10}^{\ast}}^{2}{\sigma_{s0}^{\ast}}^{2}}{(\zeta_{1}^{*}-{\psi_{1}^{*}}^{2})^{2}}+\frac{1}{\zeta_{1}^{*}-{\psi_{1}^{*}}^{2}}\bigg{)}+\\ &\qquad(1-\bar{T}){\beta_{01}^{\ast}}^{2}{\sigma_{s1}^{\ast}}^{2}\bigg{(}\frac{2{{\rho}^{\ast}}^{2}{\beta_{01}^{\ast}}^{2}{\sigma_{s1}^{\ast}}^{2}}{(\zeta_{0}^{*}-{\psi_{0}^{*}}^{2})^{2}}+\frac{1}{\zeta_{0}^{*}-{\psi_{0}^{*}}^{2}}\bigg{)}\bigg{)}^{-1},\end{split}$
@@ 4.1 Partial identification and principal causal effects
$(Y_{i},S_{i})\mid T_{i}=t,\boldsymbol{X}_{i}=\boldsymbol{x}\sim N\Bigg{(}(\mu_{yt},\mu_{sit}),\begin{pmatrix}\zeta_{t}&\psi_{t}\sigma_{st}\\ \psi_{t}\sigma_{st}&\sigma^{2}_{st},\end{pmatrix}\Bigg{)},$ (3)
where $\mu_{yt}=\lambda_{t}+\boldsymbol{\beta}_{t}^{T}(\phi_{0}+\boldsymbol{\alpha}^{T}\boldsymbol{x},\phi_{1}+\boldsymbol{\alpha}^{T}\boldsymbol{x})+\boldsymbol{\gamma}^{T}\boldsymbol{x}$, $\mu_{sit}=\phi_{t}+\boldsymbol{\alpha}^{T}\boldsymbol{x}$, $\zeta_{t}=\sigma_{y}^{2}+\boldsymbol{\beta}_{t}^{T}\boldsymbol{\Sigma}_{s}\boldsymbol{\beta}_{t}$, and $\psi_{t}=\sigma_{st}\beta_{tt}+\rho\sigma_{s,1-t}\beta_{t,1-t}$ for $t$ = 0, 1. Clearly the parameters of the observed data distribution are identifiable from the observed data, and given $\boldsymbol{X}_{i}=\boldsymbol{x}$, there are 5 identifiable parameters for each of $t=0,1$ leading to 10 identifiable parameters. Excluding the parameters corresponding to $\boldsymbol{X}$ given by $(\boldsymbol{\gamma},\boldsymbol{\alpha})$, Model (2) has 11 parameters, showing that there are fewer known equations than unknown variables. This leads to partial identification of certain parameters, despite being the simplest, parametric model possible that allows for violations of principal ignorability. Partial identification means that there are multiple points, or a region of parameter values, that are equally supported by the data and lead to the same observed data distribution, which can not be distinguished even with an infinite sample size. For a general discussion on Bayesian inference in such scenarios, we point readers to [12]. Naturally, this also implies that the parameters of more complex semi- or nonparametric models, such as those seen in the literature [4, 37], are at best partially identified in this setting unless additional constraints are imposed on the model parameters.
In the presence of such partial identification, we are left with two options. One can proceed with Bayesian inference where the parameters are only partially identified, and we detail the size of such partial identification regions in Section 4.2. Alternatively, additional assumptions can be made, which allows for point identification of the model parameters, and potentially $\rho$ as well, which we detail in Section 4.3. We also provide weaker, alternative assumptions in Section 5 that do not obtain point identification, but can drastically reduce the widths of partial identification regions.

### v2

@@ Abstract
Principal stratification is a general framework for studying causal mechanisms involving post-treatment variables. When estimating principal causal effects, the principal ignorability assumption is commonly invoked, which we study in detail in this manuscript. Our first key contribution is studying a commonly used strategy of using parametric models to jointly model the outcome and principal strata without requiring the principal ignorability assumption. We show that even if the joint distribution of principal strata is known, this strategy necessarily leads to only partial identification of causal effects, even under very simple and correctly specified outcome models. While principal ignorability leads to point identification in this setting, we discuss alternative, weaker assumptions and show how they can lead to informative partial identification regions. An additional contribution is that we provide theoretical support to strategies used in the literature for identifying association parameters that govern the joint distribution of principal strata. We prove that this is possible, but only if the principal ignorability assumption is violated. Additionally, due to partial identifiability of causal effects even when these association parameters are known, we show that these association parameters are only identifiable under strong parametric constraints. Lastly, we extend these results to more flexible semiparametric and nonparametric Bayesian models.
@@ 2.1 Notation and estimands
Throughout let $\boldsymbol{X}$ denote a $p$-dimensional vector of pre-treatment covariates, $T$ a binary treatment, $S$ a binary or continuous intermediate, and $Y$ an outcome of interest. Our observed data consists of $n$ independent copies of these random variables, and we refer to the observed data by $\mathcal{D}$. We adopt the potential outcome framework [34] and define $S(t)$ to be the potential intermediate that would be observed had treatment been set to $t$, and similarly let $Y(t)$ be the potential outcome. Throughout, we let $\boldsymbol{U}=(S(0),S(1))$ denote the joint values of the two potential intermediates, and let $S$ be the observed value of the intermediate. We assume the Stable Unit Treatment Value Assumption (SUTVA, [35]), which implies that 1) there is no ’interference,’ meaning that the potential outcomes for unit $i$ do not change with the treatments assigned to other units, and 2) there are not different versions of each treatment level. This assumption ensures that $S=S(T)$ and $Y=Y(T)$. Principal strata are defined by the joint potential values of the intermediate variable $\boldsymbol{U}=(S(0),S(1))$. The causal estimands of interest are then defined in terms of average differences in potential outcomes within principal strata, which are given by
$PCE(\boldsymbol{u})=E\{Y(1)-Y(0)\mid\boldsymbol{U}=\boldsymbol{u}\}.$
@@
$PCE(\boldsymbol{u},\boldsymbol{x})=E\{Y(1)-Y(0)\mid\boldsymbol{U}=\boldsymbol{u},\boldsymbol{X}=\boldsymbol{x}\}.$
@@ 2.2 Review of identification assumptions in PS analysis
$10<P(T=1\mid\boldsymbol{X}=\boldsymbol{x})<$ 1 for all $\boldsymbol{x}$.
@@ 2.2.1 Binary intermediates
This assumption rules out a direct effect of the treatment on the outcome that does not go through a change in the intermediate. Again, this is reasonable in randomized studies of noncompliance where randomization to a treatment should not affect the outcome, except through its impact on what treatment is actually taken. Under the exclusion restriction and monotonicity assumptions, the average causal effect for compliers is identified as the ratio of the average causal effects on $Y$ and $S$:
$E\{Y(1)-Y(0)\mid S(0)=0,S(1)=1\}=\frac{E\{Y(1)-Y(0)\}}{E\{S(1)-S(0)\}}.$
@@
$E\{Y(t)\mid S(t),S(1-t),\boldsymbol{X}\}=E\{Y(t)\mid S(t),\boldsymbol{X}\},\qquad t=0,1.$
@@
$PCE(\boldsymbol{u},\boldsymbol{x})=E(Y\mid S=s_{1},\boldsymbol{X}=\boldsymbol{x},T=1)-E(Y\mid S=s_{0},\boldsymbol{X}=\boldsymbol{x},T=0),$
@@ 2.2.2 Continuous intermediates
For a binary $S$, there are only four principal strata, and the monotonicity assumption eliminates one of these groups and enables the identification of the probability mass of $\boldsymbol{U}\mid\boldsymbol{X}$ [1]. For a continuous $S$, there are infinitely many principal strata, and while the monotonicity assumption can rule out some of them, there still remain infinitely many principal strata. Thus, monotonicity alone cannot produce the identification of the distribution of $\boldsymbol{U}\mid\boldsymbol{X}$, sometimes referred to as the principal density for continuous intermediates. Typically, researchers make parametric assumptions in order to identify this distribution. One approach is to assume the joint distribution of principal strata follows a known copula function $\mathbb{C}_{\rho}(\cdot,\cdot)$ (see [28]; [20]; [4]):
$pr(S(0)\leq s_{0},S(1)\leq s_{1}\mid\boldsymbol{X})=\mathbb{C}_{\rho}(pr(S(0)\leq s_{0}\mid\boldsymbol{X}),pr(S(1)\leq s_{1}\mid\boldsymbol{X})),$
@@
In general, the explicit identification of the correlation between potential intermediates is infeasible because only one of them is ever observed per observation. This has led to a common practice of $\rho$ being selected based on domain knowledge, and a sensitivity analysis is performed to assess the impact of this parameter. Throughout this article, we refer to parameters like $\rho$ that cannot be identified marginally as the association parameters of the principal strata. Under this distributional assumption, combined with Assumptions (1) and (5), [28] demonstrated nonparametric identification of PCEs. Given that knowledge of this correlation parameter is a strong assumption, and that results can be sensitive to its choice, many authors have taken a parametric modeling approach to estimate this correlation. Additionally, in most of these works, principal ignorability is not assumed, showing that the parametric modeling assumptions reduce the number of structural assumptions that need to be made. This is the focus of our work, as we describe the extent to which inference in these settings is possible in the following sections. While we focus primarily on continuous intermediate variables, additional discussion extending our results to binary intermediates is included in Appendix C, and, when appropriate, we will highlight differences between these two cases.
@@ 3 On the possibility of identifying $\rho$
3.1 A simple parametric model
@@
For the purpose of illustration, we first focus on a simple, linear parametric model for the outcome as well as the principal strata. There are two main reasons for considering this simple setting. For one, this is arguably the simplest modeling strategy that is making the strongest parametric assumptions, and therefore one would assume is the most likely scenario to obtain identification. However, we show that even under such strong parametric assumptions, point identification is not possible in this setting, which has implications for other approaches in the literature that utilize more complicated modeling strategies. Second, this scenario leads to simple, closed-form expressions for many partial identification regions of interest that help build intuition for the identification issues present in these models. Importantly, we extend these ideas to more complex semiparametric and nonparametric models in Section 7. We focus on a model given by
$\begin{split}E(Y_{i}\mid\boldsymbol{U}_{i},\boldsymbol{X}_{i},T_{i}=t)&=\boldsymbol{\beta}_{t}^{T}\boldsymbol{U}_{i}+\lambda_{t}+\boldsymbol{\gamma}^{T}\boldsymbol{X}_{i},\\ E(\boldsymbol{U}_{i}\mid\boldsymbol{X}_{i})&=(\phi_{0}+\boldsymbol{\alpha}^{T}\boldsymbol{X}_{i},\phi_{1}+\boldsymbol{\alpha}^{T}\boldsymbol{X}_{i})^{T},\end{split}$
where $\boldsymbol{\beta}_{t}=(\beta_{t0},\beta_{t1})$ for $t=0,1$. Our focus is on understanding when these model parameters are identified, and the extent to which they are partially identified when point identification is not possible.
As mentioned earlier, we primarily focus on settings with a continuous intermediate variable, though we present results for the binary intermediate variable setting in Appendix C. As we discuss in subsequent sections as well as in Appendix C, identification is more challenging for the continuous intermediate setting, and is therefore the focus of our manuscript. We assume Gaussian errors in the outcome model and a joint Gaussian distribution for the principal strata. We define $\mu_{yit}$ to be the conditional mean of the potential outcomes, $E[Y_{it}\mid\boldsymbol{U}_{i},\boldsymbol{X}_{i},T_{i}=t]$, and $\boldsymbol{\mu}_{si}=(\mu_{si0},\mu_{si1})$ to be the conditional mean of principal strata. Thus, a full data generating model can be written as:
@@
Of particular interest is $\rho$, the association parameter of the principal strata, which plays an important role in estimating PCEs. To the best of our knowledge, there is no literature formally studying whether it is possible to identify this association parameter, which we do in the following section.
Model (1) becomes a Bayesian model when we incorporate prior distributions for all parameters. Note that identification is inherently more nuanced for Bayesian models as informative prior distributions can lead to weak identification of model parameters. For a general discussion on Bayesian inference in such scenarios, we point readers to [12]. We do not consider that scenario here as we assume flat or uninformative priors for all parameters. When we discuss identification throughout, we are simply referring to scenarios where the likelihood is maximized at the true parameter values. Similarly, when referring to partial identification, we are focused on situations where the likelihood is maximized by a region of values that are equally supported by the observed data, which can not be distinguished even with an infinite sample size.
@@ 3.2 Identification of the association parameter of the principal strata
For a continuous intermediate, estimating average PCEs always requires estimating the principal density, which involves the association parameters. Identifying the association parameters is generally infeasible due to the lack of joint observations of principal strata. Some researchers assume these are known [21], while others have tried to estimate them using similar models to those in (1) ([37], [4]). The idea is that although $S(0)$ and $S(1)$ are never jointly observed, their association is implicitly included in the outcome model. The following result provides theoretical justification to this strategy:
@@
Suppose that $\boldsymbol{\theta}=(\boldsymbol{\beta}_{0},\boldsymbol{\beta}_{1},\lambda_{0},\lambda_{1},\sigma_{y}^{2},\boldsymbol{\gamma},\boldsymbol{\alpha},\sigma_{s0},\sigma_{s1},\phi_{0},\phi_{1})$ are known and principal ignorability fails: that is, at least one of $\beta_{01}$ or $\beta_{10}$ is nonzero. Under model (1), the posterior mode of $\rho\mid\mathcal{D}$ is consistent.
Additionally, in Appendix C we show an analogous result holds in situations with a binary intermediate. Theorem 1 shows when principal ignorability is violated, then identification of $\rho$ is indeed possible, though this result should be interpreted with caution. For one, it required the outcome model to be correctly specified with known parameters. We will see in the subsequent sections that the model parameters are only partially identified in many settings, and the implications of Theorem 1 would only be useful for estimating $\rho$ if we can consistently estimate these regression parameters. Another issue is that this result relies on principal ignorability being violated. One can show that the posterior of $\rho$ reduces to the prior distribution for $\rho$ when principal ignorability holds. This can also be seen in Theorem 2, where we show the asymptotic variance of the posterior distribution of $\rho$.
@@
$\begin{split}Var(\rho\mid\mathcal{D})&\approx n^{-1}\bigg{[}\bar{T}{\beta_{10}^{\ast}}^{2}{\sigma_{s0}^{\ast}}^{2}\bigg{\{}\frac{2{{\rho}^{\ast}}^{2}{\beta_{10}^{\ast}}^{2}{\sigma_{s0}^{\ast}}^{2}}{(\zeta_{1}^{*}-{\psi_{1}^{*}}^{2})^{2}}+\frac{1}{\zeta_{1}^{*}-{\psi_{1}^{*}}^{2}}\bigg{\}}+\\ &\qquad(1-\bar{T}){\beta_{01}^{\ast}}^{2}{\sigma_{s1}^{\ast}}^{2}\bigg{\{}\frac{2{{\rho}^{\ast}}^{2}{\beta_{01}^{\ast}}^{2}{\sigma_{s1}^{\ast}}^{2}}{(\zeta_{0}^{*}-{\psi_{0}^{*}}^{2})^{2}}+\frac{1}{\zeta_{0}^{*}-{\psi_{0}^{*}}^{2}}\bigg{\}}\bigg{]}^{-1},\end{split}$
@@ 4.1 Partial identification and principal causal effects
$(Y_{i},S_{i})\mid T_{i}=t,\boldsymbol{X}_{i}=\boldsymbol{x}\sim N\Bigg{(}(\mu_{yi},\mu_{si}),\begin{pmatrix}\zeta_{t}&\psi_{t}\sigma_{st}\\ \psi_{t}\sigma_{st}&\sigma^{2}_{st},\end{pmatrix}\Bigg{)},$ (2)
where $\mu_{yi}=\lambda_{t}+\boldsymbol{\beta}_{t}^{T}(\phi_{0}+\boldsymbol{\alpha}^{T}\boldsymbol{x},\phi_{1}+\boldsymbol{\alpha}^{T}\boldsymbol{x})+\boldsymbol{\gamma}^{T}\boldsymbol{x}$, $\mu_{si}=\phi_{t}+\boldsymbol{\alpha}^{T}\boldsymbol{x}$, $\zeta_{t}=\sigma_{y}^{2}+\boldsymbol{\beta}_{t}^{T}\boldsymbol{\Sigma}_{s}\boldsymbol{\beta}_{t}$, and $\psi_{t}=\sigma_{st}\beta_{tt}+\rho\sigma_{s,1-t}\beta_{t,1-t}$ for $t=0,1$. Clearly the parameters of the observed data distribution are identifiable from the observed data, and given $\boldsymbol{X}_{i}=\boldsymbol{x}$, there are 5 identifiable parameters for each of $t=0,1$ leading to 10 identifiable parameters. Excluding the parameters corresponding to $\boldsymbol{X}$ given by $(\boldsymbol{\gamma},\boldsymbol{\alpha})$, Model (1) has 11 parameters, showing that there are fewer known equations than unknown variables. This leads to partial identification of certain parameters, despite being the simplest, parametric model possible that allows for violations of principal ignorability. Naturally, this also implies that the parameters of more complex semi- or nonparametric models, such as those seen in the literature [4, 37], are at best partially identified in this setting unless additional constraints are imposed on the model parameters.
In the presence of such partial identification, we are left with two options. One can proceed with inference where the parameters are only partially identified, and we detail the size of such partial identification regions in Section 4.2. Alternatively, additional assumptions can be made, which allows for point identification of the model parameters, and potentially $\rho$ as well, which we detail in Section 4.3. We also provide weaker, alternative assumptions in Section 5 that do not obtain point identification, but can drastically reduce the widths of partial identification regions leading to more informative inference.